OpenStudy (anonymous):

9^x - 6 x 3^x - 27 = 0 then x is

hero (hero):

Start with x = 1 If that doesn't work, try x = 2 ..and so on.

hero (hero):

@yashkhatri

OpenStudy (anonymous):

hint assume 3^x = a and hence 9^x = (3^2)^x =(3^x)^2=a^2 the above equation reduces to a^2 -6a-27=0 first solve for a and hence for x......